OpenStudy (anonymous):

Find the center (h, k) and radius r of the circle with the given equation. x2 + y2 + 14x + 18y = -114

myininaya (myininaya):

\[x^2+14x+(\frac{14}{2})^2+y^2+18y+(\frac{18}{2})^2=-114+(\frac{14}{2})^2+(\frac{18}{2})^2\] \[(x+7)^2+(y+9)^2=-114+49+81\] \[(x+7)^2+(y+9)^2=16\] r=4 C=(-7,-9)